I have the correct answer, I just don't know how to get it??
Please help.
Correct answer. 13m + 3=6-8m
Translate the following verbal phrase into a mathematical expression using m for the variable.
3 more than the product of a number and 13 gives 6 decreased by number multiplied by 8.

Answer by mananth(16946)

3 more than the product of a number and 13 gives 6 decreased by number multiplied by 8.
let the number be m
product of a number and 13 = 13m
3more than the product
= 13m +3
..
6 decreased by a number= 6-m
multiplied by 8 means 8(6-m)
..
13m+3 = 8(6-m)
..